Maria is 3 years younger than her brother Juan the sum of their ages is 27 how old are Maria and Juan?
Answer by macston(5194)

.
J=Juan's age; M=Maria's age=J-3
.
J+M=27
J+(J-3)=27
2J=30
J=15
ANSWER 1: Juan is 15 years old.
.
M=J-3=15-3=12
ANSWER 2: Maria is 12 years old.
.
CHECK:
M+J=27
12+15=27
27=27
.